# 1S-DAug: One-Shot Data Augmentation for Robust Few-Shot Generalization

### v1

@@ 9 Impact Statement
In safety-critical deployments, any gains from augmentation should be verified with rigorous testing and human oversight consistent with regulatory standards.
Appendix A Proof of Proposition Risk Decomposition
Recall that $R(g)=\tfrac{1}{4}\,\mathbb{E}\bigl[(g(x)-y)^{2}\bigr]$ and $f,f_{A}:\mathcal{X}\to\{-1,1\}$, $\tilde{f}=\tfrac{1}{2}(f+f_{A})$. Since $f^{2}(x)=f_{A}^{2}(x)=y^{2}=1$,
$R(f)=\frac{1}{4}\,\mathbb{E}\bigl[(f-y)^{2}\bigr]=\frac{1}{4}\,\mathbb{E}[f^{2}-2fy+y^{2}]=\frac{1}{2}-\frac{1}{2}\,\mathbb{E}[f(x)y].$ (14)
@@ Appendix A Proof of Proposition Risk Decomposition
$R(\tilde{f})=\frac{1}{4}\,\mathbb{E}\Bigl[\Bigl(\frac{f+f_{A}}{2}-y\Bigr)^{2}\Bigr]=\frac{1}{16}\,\mathbb{E}\bigl[(f+f_{A})^{2}\bigr]-\frac{1}{4}\,\mathbb{E}\bigl[(f+f_{A})y\bigr]+\frac{1}{4}\,\mathbb{E}[y^{2}].$ (15)
@@
$R(\tilde{f})=\frac{1}{16}\,\mathbb{E}[2+2ff_{A}]-\frac{1}{4}\,\mathbb{E}[fy+f_{A}y]+\frac{1}{4}=\frac{1}{8}\bigl(1+\mathbb{E}[f(x)f_{A}(x)]\bigr)-\frac{1}{4}\,\mathbb{E}[f(x)y]-\frac{1}{4}\,\mathbb{E}[f_{A}(x)y]+\frac{1}{4}.$ (16)
Subtracting $R(f)$ gives
$R(\tilde{f})-R(f)=\frac{1}{4}\bigl(\mathbb{E}[f(x)y]-\mathbb{E}[f_{A}(x)y]\bigr)+\frac{1}{8}\bigl(\mathbb{E}[f(x)f_{A}(x)]-1\bigr).$ (17)
@@ Appendix B Main Theoretical Results Details
Throughout this appendix we use the pairwise reduction of Section 5.1. Each input is a query-prototype pair $x=(q,p)$ with label $y\in\{-1,1\}$. For a fixed encoder $\Phi_{\theta}$ and prototype $p$ we define the difference feature $\Omega_{\theta}(x):=\Phi_{\theta}(q)-p$ and the Euclidean score $g_{\theta}(x):=-\|\Omega_{\theta}(x)\|_{2}^{2}$. The 0-11 pairwise risk of $\theta$ is
@@
which coincides with the classification risk used in Section 5.3.
@@
B.1 Episodic Euclidean Model and Test-Time Augmentation (details)
An $N$-way $K$-shot episode has support set $S=\{(s_{c,k},c):c=1,\dots,N,\ k=1,\dots,K\}$ and query set $Q=\{(q_{j},y_{j})\}_{j=1}^{n_{q}}$ with $y_{j}\in\{1,\dots,N\}$. A single encoder $\Phi_{\theta}:\mathcal{X}\to\mathbb{R}^{d}$ produces features $z_{c,k}:=\Phi_{\theta}(s_{c,k})$ and $z_{q}:=\Phi_{\theta}(q)$. Class prototypes and the Euclidean classifier are
$p_{c}:=\frac{1}{K}\sum_{k=1}^{K}z_{c,k},\qquad\hat{y}(q):=\arg\min_{c}\|\Phi_{\theta}(q)-p_{c}\|_{2}^{2},$
where $p_{c}$ corresponds to $p_{c}$ in Section 4. For analysis we reduce episodes to binary query-prototype pairs: $x=(q,p)$ with label $y\in\{-1,1\}$, where $y=+1$ if $p$ is the prototype of the true class of $q$ and y=1y=-1 otherwise. We work with the difference feature $\Omega_{\theta}(x):=\Phi_{\theta}(q)-p$ and Euclidean score $g_{\theta}(x):=-\|\Omega_{\theta}(x)\|_{2}^{2}$, and assume a uniform pre-augmentation radius bound
$\|\Omega_{\theta}(x)\|_{2}\leq r_{0}\qquad\text{for all }x\text{ and }\theta,$
as in Section 5.1.
At test time, 1S-DAug uses support-side feature averaging and query-side logit averaging as described in Section 4, yielding (possibly aggregated) prototypes $p_{c}$. For the theory we only need query-side notation. Let $A_{k}(q)$ be the $k$-th augmented view of a query $q$ ($k=0,\dots,K_{a}$), $z^{(k)}(q):=\Phi_{\theta}(A_{k}(q))$, and $(\alpha_{k})_{k=0}^{K_{a}}$ convex weights with $\sum_{k}\alpha_{k}=1$. We define the averaged query embedding
$\bar{z}_{\mathrm{qry}}(q):=\sum_{k=0}^{K_{a}}\alpha_{k}z^{(k)}(q),$
the aggregated difference feature $\bar{\Omega}_{\theta}(x):=\bar{z}_{\mathrm{qry}}(q)-p$, and the aggregated score $\bar{g}_{\theta}(x):=-\|\bar{\Omega}_{\theta}(x)\|_{2}^{2}$, where $p$ denotes a (possibly aggregated) prototype. For squared Euclidean scores, logit averaging over per-view scores $g_{\theta}^{(k)}(q,p)=-\|z^{(k)}(q)-p\|_{2}^{2}$ is equivalent (up to a class-independent constant) to using $\bar{z}_{\mathrm{qry}}(q)$; see Appendix B.3. We therefore express all bounds in terms of $\bar{\Omega}_{\theta}(x)$ and $\bar{g}_{\theta}(x)$.
B.2 Margin-based Bound and Rademacher Complexity
For a score function $g_{\theta}$ and a pair $(x,y)\sim D$ with $y\in\{-1,1\}$, define the signed margin $u_{\theta}(x):=y\,g_{\theta}(x)$. Fix a margin parameter $\rho>0$ and the piecewise-linear margin loss
@@ B.2 Margin-based Bound and Rademacher Complexity
The (population) margin risk and empirical margin risk on a sample $S=\{(x_{i},y_{i})\}_{i=1}^{m}$ are
$R_{\rho}(\theta):=\mathbb{E}[\tau_{\rho}(u_{\theta}(x))],\qquad\widehat{R}_{S,\rho}(\theta):=\frac{1}{m}\sum_{i=1}^{m}\tau_{\rho}(y_{i}g_{\theta}(x_{i})).$
Let the empirical 0-11 pairwise risk on $S$ be
$\widehat{R}_{\mathrm{cls},S}(\theta):=\frac{1}{m}\sum_{i=1}^{m}\mathbf{1}\{y_{i}g_{\theta}(x_{i})\leq 0\}.$
@@
$R_{\mathrm{cls}}(\theta)\leq R_{\rho}(\theta)\quad\text{and}\quad\widehat{R}_{\mathrm{cls},S}(\theta)\leq\widehat{R}_{S,\rho}(\theta).$
@@
$\widehat{\mathfrak{R}}_{S}(\mathcal{G}):=\mathbb{E}_{\sigma}\Bigl[\sup_{g\in\mathcal{G}}\frac{1}{m}\sum_{i=1}^{m}\sigma_{i}g(x_{i})\Bigr],$
where $\sigma_{i}$ are independent Rademacher variables. A standard contraction argument with $\tau_{\rho}$ yields Theorem 5.2 in the main text:
@@
For any $\rho>0$ and $\delta>0$, with probability at least $1-\delta$ over $S\sim D^{m}$, every $\theta$ satisfies
$R_{\mathrm{cls}}(\theta)\;\leq\;\widehat{R}_{S,\rho}(\theta)+\frac{2}{\rho}\,\widehat{\mathfrak{R}}_{S}(\mathcal{G})+\sqrt{\frac{\log(1/\delta)}{2m}}.$
@@
Suppose each encoder $\Phi_{\theta}$ is realised by a feedforward network with 1-Lipschitz nonlinearities and layer spectral norms $\|W_{\ell}\|_{2}\leq s_{\ell}$ such that $\prod_{\ell=1}^{L}s_{\ell}\leq L_{\mathrm{enc}}$, and that $\|\Omega_{\theta}(x)\|_{2}\leq r$ for all $x$ and $\theta$. Then there exists a constant $C_{\mathrm{enc}}>0$ (depending only on the architecture) such that, for any sample $S$,
$\widehat{\mathfrak{R}}_{S}(\mathcal{G})\;\leq\;C_{\mathrm{enc}}\,L_{\mathrm{enc}}\,\frac{r}{\sqrt{m}}.$
@@
In particular, in the setting of Section 5.3, one may take $r=r_{0}$ before augmentation and $r=\hat{r}$ after augmentation.
Proof sketch.
By assumption, each encoder $\Phi_{\theta}$ is realised by a feedforward network with 1-Lipschitz nonlinearities and layer spectral norms $\|W_{\ell}\|_{2}\leq s_{\ell}$ satisfying $\prod_{\ell=1}^{L}s_{\ell}\leq L_{\mathrm{enc}}$. Standard results on spectral-norm control of deep networks (e.g. via composing linear maps and 1-Lipschitz activations) imply that $\Phi_{\theta}$ is $L_{\mathrm{enc}}$-Lipschitz with respect to the input $\ell_{2}$-norm, i.e.
$\|\Phi_{\theta}(x)-\Phi_{\theta}(x^{\prime})\|_{2}\;\leq\;L_{\mathrm{enc}}\,\|x-x^{\prime}\|_{2}\qquad\text{for all }x,x^{\prime}.$
Since $p$ does not depend on $q$ for a fixed pair $x=(q,p)$, the difference feature $\Omega_{\theta}(x)=\Phi_{\theta}(q)-p$ is also $L_{\mathrm{enc}}$-Lipschitz in $q$. On the domain where $\|\Omega_{\theta}(x)\|_{2}\leq r$ for all $x$ and $\theta$, the score $g_{\theta}(x)=-\|\Omega_{\theta}(x)\|_{2}^{2}$ is $2rL_{\mathrm{enc}}$-Lipschitz in $x$: the gradient of $g_{\theta}$ with respect to $\Omega_{\theta}$ has norm $2\|\Omega_{\theta}(x)\|_{2}\leq 2r$, and $\Omega_{\theta}$ itself is $L_{\mathrm{enc}}$-Lipschitz.
Let $\mathcal{G}=\{g_{\theta}:\theta\in\Theta\}$ be the corresponding score class. It is therefore contained in a class of Lipschitz real-valued functions whose Lipschitz constant is bounded by $2rL_{\mathrm{enc}}$. Standard covering-number/Rademacher arguments for such Lipschitz function classes (see, e.g., generic chaining bounds for $L$-Lipschitz functions on an $\ell_{2}$-ball) yield that there exists a constant $C>0$, depending only on the input dimension and hence only on the architecture, such that for any sample $S$ of size $m$,
$\widehat{\mathfrak{R}}_{S}(\mathcal{G})\;\leq\;C\,\bigl(2rL_{\mathrm{enc}}\bigr)\,\frac{1}{\sqrt{m}}.$
Setting $C_{\mathrm{enc}}:=2C$ gives
$\widehat{\mathfrak{R}}_{S}(\mathcal{G})\;\leq\;C_{\mathrm{enc}}\,L_{\mathrm{enc}}\,\frac{r}{\sqrt{m}},$
which is the claimed bound. ∎
B.3 Equivalence of Logit Averaging and Feature Averaging
For completeness we record the standard equivalence between logit averaging and feature averaging for squared Euclidean scores.
Fix a query $q$ and a prototype $p$, and let
$z^{(k)}(q):=\Phi_{\theta}\bigl(A_{k}(q)\bigr)$
denote the encoded query under augmentation $k$, with convex weights $\alpha_{k}\geq 0$, $\sum_{k}\alpha_{k}=1$. The per-view scores are
$g_{\theta}^{(k)}(q,p):=-\bigl\|z^{(k)}(q)-p\bigr\|_{2}^{2},\qquad\tilde{g}_{\theta}(q,p):=\sum_{k}\alpha_{k}\,g_{\theta}^{(k)}(q,p).$
Define the averaged query feature

### v3

@@ 9 Impact Statement
In safety-critical deployments, any gains from augmentation should be verified with rigorous testing and human oversight.
Appendix A Proof of the Risk-Decomposition Proposition
Proof.
Recall that $\mathcal{R}(g)=\tfrac{1}{4}\,\mathbb{E}\bigl[(g(x)-y)^{2}\bigr]$ and $f,f_{A}:\mathcal{X}\to\{-1,1\}$, $\tilde{f}=\tfrac{1}{2}(f+f_{A})$. Since $f^{2}(x)=f_{A}^{2}(x)=y^{2}=1$,
$\mathcal{R}(f)=\frac{1}{4}\,\mathbb{E}\bigl[(f-y)^{2}\bigr]=\frac{1}{4}\,\mathbb{E}[f^{2}-2fy+y^{2}]=\frac{1}{2}-\frac{1}{2}\,\mathbb{E}[f(x)y].$ (14)
@@ Appendix A Proof of the Risk-Decomposition Proposition
$\mathcal{R}(\tilde{f})=\frac{1}{4}\,\mathbb{E}\Bigl[\Bigl(\frac{f+f_{A}}{2}-y\Bigr)^{2}\Bigr]=\frac{1}{16}\,\mathbb{E}\bigl[(f+f_{A})^{2}\bigr]-\frac{1}{4}\,\mathbb{E}\bigl[(f+f_{A})y\bigr]+\frac{1}{4}\,\mathbb{E}[y^{2}].$ (15)
@@
$\mathcal{R}(\tilde{f})=\frac{1}{16}\,\mathbb{E}[2+2ff_{A}]-\frac{1}{4}\,\mathbb{E}[fy+f_{A}y]+\frac{1}{4}=\frac{1}{8}\bigl(1+\mathbb{E}[f(x)f_{A}(x)]\bigr)-\frac{1}{4}\,\mathbb{E}[f(x)y]-\frac{1}{4}\,\mathbb{E}[f_{A}(x)y]+\frac{1}{4}.$ (16)
Subtracting $\mathcal{R}(f)$ gives
$\mathcal{R}(\tilde{f})-\mathcal{R}(f)=\frac{1}{4}\bigl(\mathbb{E}[f(x)y]-\mathbb{E}[f_{A}(x)y]\bigr)+\frac{1}{8}\bigl(\mathbb{E}[f(x)f_{A}(x)]-1\bigr).$ (17)
@@ Appendix B Main Theoretical Results Details
Throughout this appendix we use the pairwise reduction. Each input is a query-prototype pair $x=(q,p)$ with label $y\in\{-1,1\}$. For a fixed encoder $\Phi_{\theta}$ and prototype $p$ we define the difference feature $\Omega_{\theta}(x):=\Phi_{\theta}(q)-p$ and the shifted Euclidean score $g_{\theta}(x):=\beta-\|\Omega_{\theta}(x)\|_{2}^{2}$ for a fixed decision threshold $\beta\in\mathbb{R}$. The 0-11 pairwise risk of $\theta$ is
@@
which serves as a proxy for the actual rank-based episodic classification.
@@
B.1 Episodic Euclidean Model and Test-Time Augmentation
An actual $N$-way $K$-shot episode has support set $S=\{(s_{c,k},c):c=1,\dots,N,\ k=1,\dots,K\}$ and query set $Q=\{(q_{j},y_{j})\}_{j=1}^{n_{q}}$ with $y_{j}\in\{1,\dots,N\}$. A single encoder $\Phi_{\theta}:\mathcal{X}\to\mathbb{R}^{d}$ produces features $z_{c,k}:=\Phi_{\theta}(s_{c,k})$ and $z_{q}:=\Phi_{\theta}(q)$. Class prototypes and the Euclidean classifier are
$p_{c}:=\frac{1}{K}\sum_{k=1}^{K}z_{c,k},\qquad\hat{y}(q):=\arg\min_{c}\|\Phi_{\theta}(q)-p_{c}\|_{2}^{2}.$
B.2 Equivalence of Logit Averaging and Feature Averaging
We record the equivalence between logit averaging and feature averaging.
Fix a query $q$ and a prototype $p$, and let
$z^{(\nu)}(q):=\Phi_{\theta}\bigl(A_{\nu}(q)\bigr)$
denote the encoded query under augmentation $\nu$, with convex weights $\alpha_{\nu}\geq 0$, $\sum_{\nu}\alpha_{\nu}=1$, and let $\beta\in\mathbb{R}$ be a fixed score threshold. The per-view scores are
$g_{\theta}^{(\nu)}(q,p):=\bigl\|z^{(\nu)}(q)-p\bigr\|_{2}^{2},\qquad\tilde{g}_{\theta}(q,p):=\sum_{\nu}\alpha_{\nu}\,g_{\theta}^{(\nu)}(q,p).$
Define the averaged query feature
$\bar{z}_{\mathrm{qry}}(q):=\sum_{\nu}\alpha_{\nu}\,z^{(\nu)}(q),\qquad\bar{g}_{\theta}(q,p):=\bigl\|\bar{z}_{\mathrm{qry}}(q)-p\bigr\|_{2}^{2}.$
Then
$\begin{split}\tilde{g}_{\theta}(q,p)&=\sum_{\nu}\alpha_{\nu}\bigl\|z^{(\nu)}(q)-p\bigr\|_{2}^{2}\\ &=\sum_{\nu}\alpha_{\nu}\Bigl(\|z^{(\nu)}(q)\|_{2}^{2}-2\langle z^{(\nu)}(q),p\rangle+\|p\|_{2}^{2}\Bigr)\\ &=\sum_{\nu}\alpha_{\nu}\|z^{(\nu)}(q)\|_{2}^{2}+2\Bigl\langle\bar{z}_{\mathrm{qry}}(q),p\Bigr\rangle-\|p\|_{2}^{2}\\ &=\sum_{\nu}\alpha_{\nu}\|z^{(\nu)}(q)\|_{2}^{2}-\|\bar{z}_{\mathrm{qry}}(q)\|_{2}^{2}+\|\bar{z}_{\mathrm{qry}}(q)\|_{2}^{2}+2\langle\bar{z}_{\mathrm{qry}}(q),p\rangle-\|p\|_{2}^{2}\\ &=\bigl\|\bar{z}_{\mathrm{qry}}(q)-p\bigr\|_{2}^{2}-\sum_{\nu}\alpha_{\nu}\|z^{(\nu)}(q)\|_{2}^{2}+\|\bar{z}_{\mathrm{qry}}(q)\|_{2}^{2}.\end{split}$ (18)
The last two terms depend on $q$ and the set of views $\{z^{(\nu)}(q)\}$ but not on $p$. Hence, for fixed $q$, comparing classes by $\tilde{g}_{\theta}(q,p_{c})$ is equivalent to comparing them by
$\bar{g}_{\theta}(q,p_{c})=\bigl\|\bar{z}_{\mathrm{qry}}(q)-p_{c}\bigr\|_{2}^{2}.$
In other words, query-side logit averaging with shifted squared Euclidean scores induces exactly the same class ranking as nearest-prototype classification in the feature space of the averaged query embedding $\bar{z}_{\mathrm{qry}}(q)$.
B.3 Margin-based Bound and Rademacher Complexity
Let $D$ denote the pairwise distribution of $(x,y)$ on $\mathcal{X}\times\{-1,1\}$. For a score function $g_{\theta}$ and a pair $(x,y)\sim D$ with $y\in\{-1,1\}$, define the signed margin $u_{\theta}(x):=y\,g_{\theta}(x)$. Fix a margin parameter $\rho>0$ and the piecewise-linear margin loss
@@ B.3 Margin-based Bound and Rademacher Complexity
The (population) margin risk and empirical margin risk on a generic sample $\mathcal{T}=\{(x_{i},y_{i})\}_{i=1}^{m}$ are
$R_{\rho}(\theta):=\mathbb{E}[\tau_{\rho}(u_{\theta}(x))],\qquad\widehat{R}_{\mathcal{T},\rho}(\theta):=\frac{1}{m}\sum_{i=1}^{m}\tau_{\rho}(y_{i}g_{\theta}(x_{i})).$
Let the empirical 0-11 pairwise risk on $\mathcal{T}$ be
$\widehat{R}_{\mathrm{cls},\mathcal{T}}(\theta):=\frac{1}{m}\sum_{i=1}^{m}\mathbf{1}\{y_{i}g_{\theta}(x_{i})\leq 0\}.$
@@
$R_{\mathrm{cls}}(\theta)\leq R_{\rho}(\theta)\quad\text{and}\quad\widehat{R}_{\mathrm{cls},\mathcal{T}}(\theta)\leq\widehat{R}_{\mathcal{T},\rho}(\theta).$
@@
$\widehat{\mathfrak{R}}_{\mathcal{T}}(\mathcal{G}):=\mathbb{E}_{\zeta}\Bigl[\sup_{g\in\mathcal{G}}\frac{1}{m}\sum_{i=1}^{m}\zeta_{i}g(x_{i})\Bigr],$
where $\zeta_{i}$ are independent Rademacher variables. A standard contraction argument with $\tau_{\rho}$ yields Theorem 5.2 in the main text:
@@
For any $\rho>0$ and $\delta>0$, with probability at least $1-\delta$ over $\mathcal{T}\sim D^{m}$, every $\theta$ satisfies
$R_{\mathrm{cls}}(\theta)\;\leq\;\widehat{R}_{\mathcal{T},\rho}(\theta)+\frac{2}{\rho}\,\widehat{\mathfrak{R}}_{\mathcal{T}}(\mathcal{G})+\sqrt{\frac{\log(1/\delta)}{2m}}.$
Proof.
Let
$\mathcal{G}:=\{g_{\theta}:\theta\in\Theta\},\qquad\mathcal{H}_{\rho}:=\Bigl\{h_{\theta}:(x,y)\mapsto\tau_{\rho}\!\bigl(y\,g_{\theta}(x)\bigr)\;:\;\theta\in\Theta\Bigr\}.$
We proceed in three steps.
Step 1: margin surrogate upper-bounds the classification loss.
For every $t\in\mathbb{R}$,
$\mathbf{1}\{t\leq 0\}\leq\tau_{\rho}(t).$
Indeed, if $t\leq 0$, both sides are equal to $1$; if $0<t<\rho$, then $\mathbf{1}\{t\leq 0\}=0$ while $\tau_{\rho}(t)=1-t/\rho\in(0,1)$; and if $t\geq\rho$, both sides are equal to 0. Applying this pointwise with $t=y\,g_{\theta}(x)$ and taking expectation gives
$R_{\mathrm{cls}}(\theta)\leq R_{\rho}(\theta),\qquad R_{\rho}(\theta):=\mathbb{E}_{(x,y)\sim D}\bigl[\tau_{\rho}(y\,g_{\theta}(x))\bigr].$ (19)
Step 2: uniform convergence for the surrogate class.
Each $h_{\theta}\in\mathcal{H}_{\rho}$ is $[0,1]$-valued. Hence the standard empirical Rademacher generalization bound for bounded function classes implies that, with probability at least $1-\delta$ over $\mathcal{T}\sim D^{m}$, simultaneously for all $\theta$,
$R_{\rho}(\theta)\leq\widehat{R}_{\mathcal{T},\rho}(\theta)+2\,\widehat{\mathfrak{R}}_{\mathcal{T}}(\mathcal{H}_{\rho})+\sqrt{\frac{\log(1/\delta)}{2m}}.$ (20)
Step 3: contraction by the ramp loss.
The function $\tau_{\rho}$ is $(1/\rho)$-Lipschitz on $\mathbb{R}$. Define
$\phi_{\rho}(t):=\tau_{\rho}(t)-\tau_{\rho}(0)=\tau_{\rho}(t)-1.$
Then $\phi_{\rho}$ is also $(1/\rho)$-Lipschitz and satisfies $\phi_{\rho}(0)=0$. Since the constant shift disappears inside the Rademacher average,
$\begin{split}\widehat{\mathfrak{R}}_{\mathcal{T}}(\mathcal{H}_{\rho})&=\mathbb{E}_{\zeta}\!\left[\sup_{\theta\in\Theta}\frac{1}{m}\sum_{i=1}^{m}\zeta_{i}\,\tau_{\rho}\!\bigl(y_{i}g_{\theta}(x_{i})\bigr)\right]\\ &=\mathbb{E}_{\zeta}\!\left[\sup_{\theta\in\Theta}\frac{1}{m}\sum_{i=1}^{m}\zeta_{i}\,\phi_{\rho}\!\bigl(y_{i}g_{\theta}(x_{i})\bigr)\right]\\ &\leq\frac{1}{\rho}\,\mathbb{E}_{\zeta}\!\left[\sup_{\theta\in\Theta}\frac{1}{m}\sum_{i=1}^{m}\zeta_{i}\,y_{i}\,g_{\theta}(x_{i})\right].\end{split}$ (21)
where the last step is the empirical Rademacher contraction lemma. Because $y_{i}\in\{-1,+1\}$ is fixed once $\mathcal{T}$ is fixed, the products $\zeta_{i}y_{i}$ are again i.i.d. Rademacher signs. Therefore
$\widehat{\mathfrak{R}}_{\mathcal{T}}(\mathcal{H}_{\rho})\leq\frac{1}{\rho}\,\widehat{\mathfrak{R}}_{\mathcal{T}}(\mathcal{G}).$ (22)
Combining (20) and (22), we obtain
$R_{\rho}(\theta)\leq\widehat{R}_{\mathcal{T},\rho}(\theta)+\frac{2}{\rho}\,\widehat{\mathfrak{R}}_{\mathcal{T}}(\mathcal{G})+\sqrt{\frac{\log(1/\delta)}{2m}}.$
Finally, combining this inequality with (19) gives
$R_{\mathrm{cls}}(\theta)\leq\widehat{R}_{\mathcal{T},\rho}(\theta)+\frac{2}{\rho}\,\widehat{\mathfrak{R}}_{\mathcal{T}}(\mathcal{G})+\sqrt{\frac{\log(1/\delta)}{2m}},$
as claimed. ∎
@@
Suppose each encoder $\Phi_{\theta}$ is realised by a feedforward network with 1-Lipschitz nonlinearities and layer spectral norms $\|W_{\ell}\|_{2}\leq s_{\ell}$ such that $\prod_{\ell=1}^{L}s_{\ell}\leq L_{\mathrm{enc}}$, assume $B_{X}:=\sup_{q\in\mathcal{X}}\|q\|_{2}<\infty$, assume for simplicity that the prototypes are fixed, and assume that $\|\Omega_{\theta}(x)\|_{2}\leq r$ for all $x$ and $\theta$. Then there exists a constant $C_{\mathrm{enc}}>0$ (depending only on the architecture) such that, for any sample $\mathcal{T}$,
$\widehat{\mathfrak{R}}_{\mathcal{T}}(\mathcal{G})\;\leq\;C_{\mathrm{enc}}\,B_{X}\,L_{\mathrm{enc}}\,\frac{r}{\sqrt{m}}.$
@@
In particular, in the setting of Section 5.3, one may take $r=r_{0}$ for the original score class before augmentation, and for the aggregated score class one may take $r=\hat{r}$ after augmentation.
Proof.
Fix a generic sample $\mathcal{T}=\{x_{i}\}_{i=1}^{m}$ with $x_{i}=(q_{i},p_{i})$, and define the vector-valued difference-feature class
$\mathcal{F}:=\bigl\{x\mapsto\Omega_{\theta}(x)=\Phi_{\theta}(q)-p\;:\;\theta\in\Theta\bigr\}\subseteq\mathbb{R}^{d},$
as well as the vector-valued encoder class on queries